# Extreme heat is associated with reductions in human activity

### v1

@@ Abstract
Extreme heat is a growing threat to both individual livelihoods and broader economies, killing an increasing number of people each year as temperatures increase and limiting productivity—often in the countries that would benefit most from economic growth. Many studies document the link between heat waves and mortality or morbidity, and others explore the economic consequences of them, but few are able to identify the ways in which populations respond to the shock of extreme heat. Toward this end, we investigate the link between human mobility and fluctuations in temperature. Using additive models fitting data in Indonesia, India and Mexico, we show that extreme heat reduces mobility by up to 10% in rural areas and 5% in urban settings. Effects are stronger in poorer areas. Twinning these models with climate projections, we show that without adaptation mobility will fall 1-2% per year on aggregate, with certain seasons and places seeing activity fall by as much as 10%. According to our estimates, rural areas will face the highest relative losses and urban megacities will experience the greatest absolute impacts.
@@ Introduction
Extreme heat threatens lives annually across the world [1, 2]. Various studies establish a relationship between rising temperatures and increases in death rates [3] and hospital admissions [4]. The worst effects of high temperatures occur in populations that are unable to cope with its effects, including the very young and the very old [1, 5]. In addition to these human costs, heat waves bring substantial economic costs, reducing productivity [6] and slowing growth [7]. This problem will likely get worse: heat waves are increasing in both magnitude and frequency [8], with duration and intensity growing since 1950 [9], and increasing at an increasing rate since 1980. Yet literature exploring the behavioral aspects of heat, which underpin these human and economic costs, has been limited in scope and scale. Here we address this gap by joining mass mobility data across 3 developing countries—India, Indonesia and Mexico—with a combined 1.8 billion people across 8 climate zones [10] with data on temperature, precipitation and solar radiation to understand how extreme heat changes activity patterns in both rural and urban contexts.
However, the literature on extreme heat shows that the impact of heat waves is not uniform across all populations. Differences arise due to two primary factors: where a person lives and who the person is. Regarding the latter, age, gender, and specific physical conditions can significantly influence how individuals are affected by extreme heat [11, 12]. As for the former, some cities—or even parts of cities—are more resilient to heat waves than others due to various factors such as proximity to water [13], urban design [14], and other geographic characteristics [15]. Additionally, in areas where people can afford cars, air conditioning, or other amenities, the impact of heat is less significant [16]. Consequently, heat waves tend to disproportionately affect different socioeconomic groups, which is evident when examining changes in their mobility patterns. For instance, areas with large numbers of disadvantaged social groups experience more severe disruptions in mobility due to temperature changes [16].
Data from the US give a window into the relationship between extreme heat and wealth. The association between heat and mortality has fallen in the US since the diffusion of air conditioning, an indication that wealth alleviates the worst effects of climate change [17]. Evidence suggests that quality of life is highest for Americans at average daily temperatures of $18^{\circ}$, and that a given person would pay more to avoid heat than cold [18]. Time use data indicate that individuals exposed to high temperatures at work reduce hours worked and that outdoor leisure is redirected indoors [19]. While this context is considerably different to that of our study here, these results serve as indications for how affluent groups—who can afford to adapt both environments and behaviors—respond to the stresses of heat.
Differences between socio-economic groups are evident not only in their initial reactions to extreme weather events but also in their recovery dynamics. With the availability of granular data capturing human movement, it has been observed that although wealthier groups initially responded more quickly to certain weather events, residents from poorer areas returned to their prior mobility patterns faster [20]. This may be in part because wealthier individuals have the means to reduce their mobility for longer periods. However, more importantly, it suggests that poorer groups face a higher risk of exposure to extreme weather, as they may be compelled to resume activity sooner, despite hazardous conditions.
Yet mortality decreases as the frequency of high temperatures increases, suggesting that populations are able to compensate or adapt to high, albeit frequent, temperatures [21]. Further, the most frequent temperature in an area is often the temperature at which the fewest deaths occur [22]. The possibility for compensatory behavior has led governments and researchers to classify extreme heat alongside hurricanes and cyclones as disruptive events, going so far as to name heat waves as we do storms [23] and develop “heat action plans” [24] to address the problem.
Extreme heat is associated with falling productivity in manufacturing [6] and agriculture [25]; it reduces economic growth in the poor countries that would benefit most from it [7]. These effects are not limited to durable changes in climate: there is evidence that anomalous hot days, rather than long periods of hot temperatures, are enough to curb productivity [26]. A growing body of evidence also indicates that heat waves cause a variety of other social ills, from increasing conflict [27] to decreasing learning amongst pupils in schools without air conditioning [28]. With further implications for the economy, heat waves cause problems for both the built and natural environments, stressing railways and roadways [29], as well as trees and plants [30].
The mechanisms by which heat influences economic behavior are clear: stress from extreme heat causes dehydration and fatigue [31], with particular risk to the elderly. For professions demanding physical labor, heat stress and heat stroke may cause organ damage or death, and can exacerbate existing health problems [24]. In addition to its impacts on physical ability, heat waves also impact cognitive ability and focus [32]. There is evidence that heat reduces hours worked and limits productivity [33]. It primarily affects agricultural labor but it also affects construction and manufacturing [34].
Research, primarily in the US but of particular relevance to this study, finds a connection between heat waves and urban activity. Rising temperatures can alter the types of activities people choose to engage in, the modes of transport they opt for, the distances they are willing to travel, and the duration of their outings [35, 36]. Notably, research indicates that during periods of heightened temperature, people tend to avoid short trips [16] and visit different types of places [37], suggesting that certain trips are not worth the trouble during extreme heat. They also use public transport less [38] and walk less [39], suggesting that comfort is a factor.
Here we attempt to isolate the behavioral foundations of these economic and human consequences by exploring how human mobility responds to changes in temperature. Research into the behavioral implications of extreme heat is growing in importance as heat waves become frequent [40]. Although studies have been able to identify economic consequences of high temperatures, the behavioral foundations of these consequences—how human behavior responds to heat, producing these effects—remains elusive. The same holds for mortality and morbidity: we know that extreme heat kills, but we know less about how individuals or groups of individuals change planned activities to reduce the likelihood of harm.
Operating under the hypothesis that extreme heat curbs mobility, the following paper asks how extreme heat changes mobility. We explore these behavioral foundations of mortality, morbidity and productivity with mass GPS signal data from mobile phones across 3 countries. We focus on India, Indonesia and Mexico—countries with large populations and different climates. Studying countries that are already hot lends importance to our study: in economic simulations, the coldest parts of the world see productivity rise with temperatures while the hottest parts experience the opposite effect [41]. We further disaggregate our analysis according to population and deprivation in order to understand how different groups in these countries respond to and recover from heat waves. Our approach sheds new light on how people manage extreme heat in developing countries and how socio-economic conditions can help or harm these efforts, and it provides the first attempt to model these behaviors across multiple climates and cultures.
@@ Results
We use daily flows between origin and destination geohash5 aggregations, comprising $\sim 236$ million trips and $\sim 1.4$ billion unique GPS signals, gathered from mobile devices in the course of daily usage (for descriptive statistics and illustrations of the network, see Fig. S1 in the supplemental information). For each geohash5 aggregation, we acquire daily temperature, precipitation and solar radiation data from the ERA5 program [42]—data resolved enough that we can observe the temperature gradient within a region, as well as across the country. Joining the two datasets, we are able to make inferences about the effect of temperature on flows to, from and between geohashes.
To begin, we use a binary classification of extreme heat to establish a connection between temperature and flows, borrowing from earlier work the definition of a heat wave as at least 3 days of top decile temperature for a region [43]. We identify periods that meet this criterion in 2019 in India and Indonesia in Fig. 1A. We use spatial interaction networks for Delhi and Jakarta, where each node is a geohash5 and each edge is the average daily trips between those nodes. We calculate these averages for the period of the heat wave and then for the periods immediately preceding and following it, comparing the within and without heat wave values. The results are shown in Fig. 1B and D: on average, edge weights (trips between geohashes) fell by 20% in Delhi and 30% in Jakarta; in the former there are some edges with gains and others with losses while in the latter the drop in activity is uniform. We use cities that did not experience heat waves as a comparison, shown in Fig. 1C, noting that much of the difference in Jakarta relative to Medan, is the result of origin-destination pairs with some trips then having 0 during the heat wave; Delhi saw a broader shift relative to Mumbai. These experiments, however, are only an initial attempt to understand the relationship between temperature and activity, and do not establish a causal link; next we model the relationship formally.
Figure 1: Extreme heat and mobility. A Temperatures in India and Indonesia throughout 2019, highlighting the hottest months of the year for both countries. B and D Change in mean daily visitation in Delhi and Jakarta between hottest week of the year and the weeks preceding and following it; activity levels fell 30% in Jakarta and 20% in Delhi. The results may be more mixed for Delhi because the hottest week of the year came in the middle of a 5-week heat wave which had already begun to influence activity patterns. C The change in distributions for Delhi and Jakarta, along with control cities that did not experience heat waves: while Delhi sees a broad shift relative to Mumbai, the lower average in Jakarta is driven by origin-destination pairs that fall from some trips to 0 during the heat wave.
Modeling the effect of extreme heat on mobility
Because the network tends to be sparse and our concern in this study is “activity”, we can define it as the number of journeys terminating in a given geohash5. In order to understand the effect of incremental changes in temperature, we model the relationship between temperature and activity using generalized additive models (GAMs), which combine the flexibility of non-parametric models with the interpretability of parametric approaches. GAMs use smooth functions that allow coefficients to vary across different points in the function, and they are appropriate for our study in particular because of the documented nonlinear effects of temperature on economic indicators like productivity [44]. (See Methods section for more detail.) We use a directed acyclic graph (DAG) to identify the minimally sufficient adjustment set—the controls that allow us to infer causal links between our independent variable of interest (temperature) with our dependent variable (activity). Shown in Fig. 2A, we identify temporal and spatial factors (identified with blue), like seasonality, deprivation and population, as determinants of activity, along with other aspects of weather—solar radiation and precipitation. Spatial fixed effects at the level of the geohash5 close all geographic confounds and further controls for weather and seasonality close the rest. In alternative specifications, we relax these fixed effects to allow for the exploration of additional factors like population and deprivation (shown in red).
Our first model considers the effect of temperature fluctuations, relative to the local average. Because prior research indicates that frequent temperatures are less lethal [21, 22], we expect that it is these relative changes rather than absolute values that matter. Fig. 1A and B show the relationship between changes in temperature (by standard deviation) and trip counts and durations (measured in minutes), respectively. Our results agree with existing literature: as temperature increases relative to the local average, activity according to these 2 measures decreases. For trip counts, the results are mixed for decreases in temperature, with strong negative effects in Indonesia and weak positive effects in India.
@@
We see in Fig. 2B that deviation from the mean temperature—in either direction—in a location is associated with reductions in activity. Showing that relative changes in temperature are better predictors than absolute changes, our results are consistent with the literature that suggests climate preferences are set locally [18]. This also agrees with research that finds that the most frequent temperature in an area is often the least fatal for that area’s population [22]. Exploring other possible changes to activity, Fig. 2C shows the partial effect of temperature anomaly on trip length: while the number of trips fall during extreme high and low temperatures, the lengths of those trips also fall.
@@
We then look at absolute temperatures to understand if levels are also a factor in driving activity. Because prolonged heat waves could elicit changes in behavior relative to shorter ones, we test our model using rolling averages of different lengths—because higher rolling averages indicate stronger or longer fluctuations in temperature. Fig. 3A shows that in India, when the temperature exceeds $33^{\circ}$, the effect of temperature on activity becomes negative; Fig. 3B shows that in Indonesia, this change from positive to negative occurs at $35^{\circ}$. A key observation for India is that during persistent heat waves, captured with longer rolling averages above $42^{\circ}$, the model projects relative adaptation, with the effect size approaching zero. While inconclusive, this suggests that extended periods of high temperatures necessitate a return-to-activity, or that policy and behavior adapt only after multiple days of extreme heat.
Finally, we are interested in the ways different demographic and geographic factors influence activity. This requires us to relax our fixed effects so that we can control for variance along other dimensions: when we control for minimum spatial level of aggregation, spatial attributes cannot vary within that level. In this specification, we use geohash3 fixed effects but add controls for population [45] and deprivation [46] at the level of geohash5. This model accounts for 57% of the variance, compared to 60%, suggesting we reclaim much of the benefit of earlier models while allowing us to explore further relationships. In order to understand the interaction between temperature and deprivation, we explore the space of model predictions across all combinations of temperature and deprivation. These predictions are shown in Fig. 3C and D, for India and Indonesia, respectively.
@@
In India, we recover the same arc as above, where activity partially recovers at the highest extremes. The contours in 3C are flatter for low deprivation and steeper for high deprivation, indicating that much of the worst effects of heat in India are felt by the poorest. Looking at Indonesia in 3D, higher temperatures bring lower activity for all levels of deprivation, but the steeper contours at higher levels of deprivation indicate that the effect is stronger for poor areas than for wealthy ones.
@@
Figure 3: The effect of extreme heat on activity. A Focusing on India, given the propensity for heat waves in India, we measure the duration and intensity of extreme heat by calculating a rolling average, finding that higher temperatures for longer are associated with stronger effects between $35-40^{\circ}$ but not above that, which may indicate compensatory behavior at a certain point. B We fit the same model for Indonesia and find similar results, with an ideal temperature and declining activity when heat surpasses it (noting that Indonesia does not have the same extremes, as it is situated within the tropics). Lasting heat waves, measured in rolling averages, show less impact in Indonesia. C and D Looking at the interaction between deprivation and temperature in India and Indonesia, respectively, we see a curve that follows earlier modeling for all socio-economic strata but it is steepest, indicating the strongest effects, for the poorest strata.
@@
As a robustness check, we use an ARIMAX model to fit the time series of the data for each geohash3 (because the geohash5 level often has incomplete data across the year). This model predicts activity by considering activity at time $t$ using activity at $t-1$ and $t-7$ along with linear and nonlinear terms for temperature, which allows the relationship to curve as temperature increases. This method implicitly controls for seasonality and day-of-week effects by using the previous day and the same day of the previous week to make a prediction, and achieves a good fit, shown for India in Fig. 4A. Important to this study, the quadratic term is negative across most regions, indicating that at a certain point the effect of temperature on activity turns negative. Fig. 4B shows, again for India, the sign and magnitude for each geohash3, which shows mixed results but less of temperature effect in the South, where temperatures are more consistent, and more of an effect in the North and Northwest. Some coefficients are positive with high magnitude, but these tend to be in the mountainous North, which does not have as many hot days, and the tropical South, which has less variability in temperature. We take the average coefficient for $\text{Temperature}^{2}$ across all countries in Fig. 4C. Note that we need to rescale the trip distribution in each country for the effect sizes to be comparable, so the value of each coefficient is difficult to interpret (in Fig. S2, we plot the curves from each ARIMAX formula to show how $\text{Temperature}^{2}$ influences the prediction). We filter out geohashes according to the length of the time series, because very short series could bias the results; generally, as this threshold of minimum observations rises, so too does the effect size.
@@
Finally, we identify the optimal temperature by making predictions across the spectrum for each geohash3 and observing where the predictions begin to fall. According to these predictions, the mean and median optimal temperatures for India are $31.1^{\circ}$ and $31.8^{\circ}$—in line with our estimates from the GAM. (These mirror what we see in S2, which plots the prediction curves for each geohash3 ARIMAX.) Although not a perfect match with the above findings, these results strengthen our assertion that extreme heat can limit activity. We document a noisy relationship between temperature optima across geohashes and the mean temperature of those geohashes ($\rho=0.25$ across all countries, $p<0.05$) in Fig. 4D: as mean temperature increases, the optimal temperature for activity also increases. This provides further evidence that populations adapt to climatic conditions.
@@
We next turn to models of Earth’s climate [47] to make estimates of how activity will be impacted in future. In particular we use data from CMIP6 simulations, extracting the estimates from the Max Planck Institute because these estimates show the best fit to past trends [48]. We then simply change the temperature-day value in 2019 with temperature-day value in 2050 according to these models. We make these projections knowing that our own study suggests that optimal temperatures are often related to the climate of the area, and thus that the population of a given area may be able to adapt to local temperatures. Nevertheless, our data suggest that optimal temperature is not always average temperature, indicating that rapid deviation from these temperatures as global warming accelerates [9] might still be a shock to communities who have acclimated to past temperatures. Furthermore, our data suggest that poor areas suffer the worst effects, and these areas might be the least able to cope with changes to the local climate going forward. As a modeling exercise, this provides us with a worst-case-scenario, but a plausible one at that.
@@
According to CMIP6 estimates shown in Fig. 5A we can expect 100-150 days per year across Northern states like Rajasthan to exceed $40^{\circ}$ and those same estimates B suggests 25 days per year will exceed $45^{\circ}$. On any given day in June when such temperatures are most likely, not a weekend nor a holiday, our model estimates that a single day at $40^{\circ}$ temperatures to reduce activity in urban areas by 2% and in rural areas by 6%, shown in Fig. 5C. Because cities generate more activity, they experience a small relative change but a large absolute loss. As we see in Fig. 5D, these losses grow to 4% and 8% for urban and rural areas, respectively, during a single day of temperatures at $40^{\circ}$.
CMIP6 estimates mapped in Fig. 6A show that many days per year will exceed $35^{\circ}$ in Sumatra and Borneo. While temperatures of $40^{\circ}$ will be rarer, shown in B they are forecast to happen despite never appearing in our 2019 data. Because Indonesia is in the tropics and does not experience much temperature variability, our model projects strong effects for heat that exists beyond the current distribution. Indonesia is planning to move its capital to the island of Borneo, from Java, by 2045. While this megaproject intends to curb the effects of subsidence, temperatures in this new region, according to CMIP6 data, will experience more hot days than in the existing capital region. Shown in Fig. 6C and D, the consequences of these temperatures, according to our modeling, are similar to the effects we see in India: rural areas with larger relative effects and urban areas with larger aggregate disruptions.
@@
In the supplemental information (Fig. S3), we show the same results for Mexico, which show the same strong effects from extreme heat and mild effects from temperatures that will occur often. An important difference to note is that while the Southern, tropical areas in Mexico will have the highest average temperature, with Fig.LABEL:S3A showing most days over $35^{\circ}$ by 2050, but the desert near the Northern border—where manufacturing serves the US market—will experience more heat waves, highlighted in Fig. LABEL:S3B. Because the literature indicates some adaptation to heat, these heat waves in the North could disrupt activity more than stably higher temperatures in the South.

### v2

# Remote sensing and GPS mobility reveal heat’s impact on human activity across diverse climates

@@ Abstract
Extreme heat is a growing threat to both individual livelihoods and broader economies, killing a growing number of people each year as temperatures rise in many parts of the world and limiting productivity. Many studies document the link between heat waves and mortality or morbidity, and others explore the economic consequences of them, but few are able to determine how populations respond to the shock of extreme heat in day-to-day activity. Toward this end, we investigate the link between human mobility and ambient temperature. Examining Indonesia, India and Mexico, we show that extreme heat reduces mobility by up to 10% in urban settings, with losses concentrated midday. We examine the shape of the relationship, finding that while heat reduces activity, very hot days and very long heat waves may induce more of it, indicating different adaptation. Effects are stronger in poorer areas. Twinning these models with climate projections, we show that without adaptation mobility may fall 1-2% per year on aggregate, with certain seasons and places seeing activity fall by as much as 10%. According to our estimates, small cities will face the highest relative losses and large cities will experience the greatest absolute impacts.
@@ Introduction
Extreme heat threatens lives globally [1, 2], with rising temperatures linked to greater mortality [3] and morbidity [4]. Vulnerable groups, including the very young and old, suffer the worst outcomes [1, 5]. Beyond human health, heat waves induce economic damage by lowering productivity [6] and slowing growth [7]. The severity of this challenge continues to mount: heat waves are intensifying in both magnitude and frequency [8], with their duration growing since 1950 and that rate of growth accelerating since 1980 [9]. Yet, the behavioral dimensions of heat impacts are not clear and research is often limited by context, often focusing on advanced economies [10]. By pairing weather and socio-economic data with GPS mobility, we show that extreme heat fundamentally disrupts activity urban areas. In doing so, we demonstrate a scalable approach for modeling human adaptation to extreme heat across diverse contexts.
The consequences of extreme heat differ based on who people are and where they live. Age, gender, and surrounding conditions shape vulnerability [11, 12], while geographical features like water bodies [13], urban design [14], and broader geographic characteristics [15] alter heat resilience. Socioeconomic status also matters: in areas with abundant air conditioning or car ownership, heat waves pose fewer disruptions [16], whereas disadvantaged communities experience sharper mobility shifts due to hot weather [16]. In the US, wealth reduces heat-related mortality via air conditioning [17], as people pay more to avoid heat than cold [18], cut working hours, and redirect outdoor leisure indoors [19]. Moreover, although wealthier populations can better delay activities, poorer ones resume hazardous routines sooner, facing heightened exposure [20].
Populations do adapt over time: mortality decreases with more frequent high temperatures, implying compensatory mechanisms [21], and the least deadly temperatures often match local climate norms [22]. Governments now classify extreme heat alongside hurricanes and cyclones, going so far as to name heat waves [23] and devise “heat action plans” [24]. Yet adverse outcomes persist: extreme heat lowers productivity in manufacturing [6] and agriculture [25], dampens growth in poorer regions [7], and a single anomalously hot day can curb output [26]. Heat also contributes to conflict [27], erodes learning in non–air-conditioned classrooms [28], and strains infrastructural [29] as well as natural systems [30]. Consequently, understanding how people adjust their daily activities is crucial.
Heat exerts its effects via physiological stress—causing dehydration, fatigue, and organ damage, especially among older individuals [31, 24]—and by compromising cognitive performance [32]. This leads to fewer working hours [33], constrained productivity in physically demanding sectors such as construction [34], and altered urban behavior. Rising temperatures reshape activity patterns in cities worldwide [35, 36], driving individuals to forgo short trips [16], shift destination choices [37], and reduce both public transport use [38] and walking [39]. These behavioral modifications, rooted in discomfort or safety concerns, underlie broader economic and social impacts. As climate change intensifies, deeper knowledge of how heat influences collective movement has become a research priority [40].
In this paper, we investigate how extreme heat modifies urban mobility in three developing countries—India, Indonesia, and Mexico—home to 1.8 billion people and spanning eight climate zones [41]. We combine mass GPS-based mobility data with thermal comfort measures to quantify the form and magnitude of heat’s influence in 2019. By modeling changes in trips, distances, and destinations, we illuminate the behavioral foundations of adverse health and economic outcomes. Given that hotter regions suffer the greatest productivity losses under climate change [42], understanding these responses is essential. We find that extreme temperatures significantly alter mobility patterns, with activity levels dropping by up to 10% during heat waves. We further disaggregate our analysis by population density and socio-economic status, yielding comparative insights into how different groups manage and recover from extreme heat across diverse climates and cultures, showing that economically disadvantaged areas show stronger reductions in mobility. While populations attempt to compensate by shifting activities to cooler hours, this adaptation is less effective in poorer areas, likely due to limited access to cooling infrastructure and inflexible work schedules. Our findings reveal how climate change could exacerbate existing inequalities in developing countries, where the most vulnerable populations face the greatest disruption to their daily activities. Since urban economies depend on the mobility and interaction [43], our results suggest that without interventions to help populations adapt to extreme heat, rising temperatures will impose a cost that falls most on those who can least afford it.
@@ Results
Impact of hot days on mobility
We use daily flows between origin and destination geohash5 aggregations, comprising $\sim 236$ million trips and $\sim 1.4$ billion unique GPS signals, gathered from mobile devices in the course of daily usage (for descriptive statistics and illustrations of the network, see Fig. S1 and S2 in the supplemental information). As these data best represent urban populations, we remove rural areas—clipping each network to “functional urban areas” [44]; Because the network is sparse and our concern is “activity”, we define it as the number of journeys ending in each geohash5. For each geohash5, we acquire Universal Thermal Climate Index (UTCI) [45] from ERA5-HEAT [46], which takes wind, sun, moisture, and temperature into account to produce a measure of how hot or cold it feels. Joining the datasets, we are able to make inferences about the effect of temperature on moves terminating in a given geohash5. (See Methods for greater detail.)
We show temperature fluctuations across India, Mexico and Indonesia throughout the year in 1A, finding that each country gives us insights into different climates: located on the equator, Indonesia has a stable climate; Mexico and India have more pronounced curves with distinct profiles. Mexico City suffered heat wave in July 2019 (3+ days in the highest temperature decile [47]), and we show in 1B mean activity per grid cell during the heat wave compared to the weeks before and average, for comparison. Although noisy, activity falls in the center, with mixed effects in the periphery. With preliminary evidence that heat changes behavior, we use a TWFE approach to understand the effects across all countries systematically. Because frequent temperatures are less lethal [21, 22], we examine deviation from the mean temperature in a geohash5. The results are shown in 1C, which plots the change relative to the strength of temperature anomaly. The results are clear in Mexico: higher temperatures mean fewer trips in a cell. In India, heat reduces afternoon activity—when temperatures peak—but activity shifts to the evening. Sensibly, the effect is inverted for cold, where waiting for the air to warm is better than during extreme heat. In Indonesia, there is limited variation and the effects we can observe are de minimis, save for the possibility—given large standard errors—that cold appears to decrease activity throughout the day while heat appears to increase it.
Figure 1: Extreme heat and mobility. A Temperatures in India, Mexico and Indonesia throughout 2019, showing that we have three different climate profiles: equatorial—and thus stable—in Indonesia, and tropical but varied and seasonal in India and Mexico. B Change in mean daily visitation in Mexico City between a heat wave in July and the weeks preceding and following it; activity levels fell 10% on aggregate, and although some peripheral areas gained foot traffic, losses are concentrated in the central city. C Results from a TWFE regression showing that, with controls for location and date, activity falls most in the afternoon on hot days—when the temperature is hottest—in Mexico and India. In Mexico activity falls throughout the day, while in India activity rises in the evening, after it has cooled off, on hotter days. Activity rises in the afternoon on cold days, suggesting that on cold days, the warmest hours of the day are conducive to activity, while on hot days, the hottest hours are destructive.
Insights into the shape of the temperature-mobility relationship
To capture incremental temperature changes, we model the temperature–activity relationship using a generalized additive model (GAM), which offers non-parametric flexibility and parametric interpretability. Our GAM uses smooth functions that accommodate varying coefficients and suits the documented nonlinear temperature effects on productivity [48]. (See Methods.) We use a directed acyclic graph (DAG) to identify the minimally sufficient adjustment set—the controls that allow us to infer causal links between our independent variable of interest (temperature) with our dependent variable (activity). For controls, we identify temporal and spatial factors as determinants of activity and use geohash5 fixed effects along with splines to capture trends over time. In alternative specifications, we interact temperature with population and deprivation to explore how these moderate its effects. While the GAM captures the shape of relationship, we use another TWFE specification with various temperature bins to provide more rigorous identification by accounting for both location-specific characteristics and time-varying factors.
We compare both models in Fig. 2, finding consistent results. In Fig. 2A we see that in India, activity declines during the afternoon but recovers in the evenings. Notably, afternoon temperature peaks have inverse effects: most active on cold days, when warmth is beneficial, but least active on hot days. In Mexico, shown in 2B, higher temperatures are associated with fewer trips during all periods of the day. We see limited effects in Indonesia in Fig. 2C, with warmer temperatures increasing activity and cooler temperatures decreasing activity—though we cannot reject the null hypothesis here.
Generally, deviation from a location’s mean temperature—hot or cold—reduces activity. Showing that relative changes in temperature are better predictors than absolute changes, our results are consistent with the literature that suggests climate preferences are set locally [18]. This also agrees with research that finds that the most frequent temperature in an area is often the least fatal for that area’s population [22].
We also explore daily effects in the supplemental information. Again, we expect that it is relative changes rather than absolute values that matter, so we consider deviations from the average. Fig. S3B and C show the relationship between changes in temperature and trip counts and durations (measured in minutes) for full days. Our results agree with existing literature: as temperature increases relative to the local average, the number of trips falls. We also conduct placebo tests in S4 and sensitivity analysis in S5 to show that our results are unlikely to be spurious.
@@
We then look at absolute temperatures to understand if levels are also a factor in driving activity. To see if prolonged heat waves differ, we test rolling averages of different lengths, where higher averages indicate stronger or longer fluctuations. Fig. 3A shows that in India, when the temperature exceeds $35^{\circ}$, the effect of temperature on activity becomes negative; Fig. 3B shows that in Mexico, this change from positive to negative occurs at $30^{\circ}$. In India, persistent heat waves (rolling averages above $45^{\circ}$) show relative adaptation, with effect sizes rising above zero. While not conclusive, this suggests that extended periods of high temperatures necessitate a return-to-activity, or that policy and behavior adapt only after multiple days of extreme heat. Replicating what we saw in the TWFE analysis, when we partition the data by time of day in Fig. S6, we see that the effect converging toward 0 at $42^{\circ}$ in India at high temperatures is the product of countervailing forces: extreme heat increases activity in the evenings while decreasing it in afternoons, when the temperature is highest. This difference is only present in India, where temperature is more variable, but it indicates that compensatory changes are at play at the extremes. In Mexico, all times of day see reductions at high temperatures.
Heterogeneities in the impact of heat on mobility
We next examine how demographic and geographic factors shape activity. In this specification, we use geohash5 fixed effects but add interaction terms for deprivation [49] or population [50]. In order to understand the interaction between temperature and deprivation, in particular, we explore the space of model predictions across all combinations of temperature and deprivation. These predictions are shown in Fig. 3C and D, for India and Mexico, respectively. (We use the same strategy for population in S6.)
@@
In India, we recover the same arc as above, where activity partially recovers at the highest extremes. The contours in 3C are flatter for low deprivation and steeper for high deprivation, indicating that much of the worst effects of heat in India are felt by the poorest. Looking at Mexico in 3D, higher temperatures bring lower activity for all levels of deprivation, but the contours at higher levels of deprivation indicate that the effect is stronger for poor areas than for wealthy ones. We see recovery at extreme heat in India but not in Mexico, where contours steepen at higher temperatures. This has important implications for vulnerable populations, as we might expect competing forces to be at play: the poor are also less able to afford taking time off from work, but they may be less able to access cooling, which could suppress activity. We see evidence that the latter pressure dominates the former, and activity is suppressed at extremes.
@@
Figure 3: The effect of extreme heat on activity. A Focusing on India, given that country’s propensity for heat waves, we measure the duration and intensity of extreme heat by calculating a rolling average, finding that higher temperatures for longer are associated with stronger effects between $40-45^{\circ}$ but not above that, which may indicate compensatory behavior at a certain point. B We fit the same model for Mexico and find similar results, with an ideal temperature and declining activity when heat surpasses it (noting that Mexico does not have the same extremes). Lasting heat waves, measured in rolling averages, show a stronger impact in Mexico and the rebound we see in India never becomes positive. C and D Looking at the interaction between deprivation and temperature in India and Mexico, respectively, we see a curve that follows above modeling for all socio-economic strata but it is steepest, indicating the strongest effects, for the poorest strata.
@@
As a robustness check and to explore spatial variation, we use an ARIMAX model at the geohash3 level (to ensure full time series data for each). This model predicts activity by considering activity at time $t$ using activity at $t-1$ and $t-7$ along with linear and nonlinear terms for temperature, which allows the relationship to curve as temperature increases. This method implicitly controls for seasonality and day-of-week effects by using the previous day and the same day of the previous week to make a prediction, and achieves a good fit, shown for India in Fig. 4A. Important to this study, the quadratic term is negative across most regions, indicating that at a certain point the effect of temperature on activity turns negative. Fig. 4B shows, again for India, the sign and magnitude for each geohash3, which shows mixed results but less of temperature effect in the South, where temperatures are more consistent, and more of an effect in the North and Northwest. Some coefficients are positive with high magnitude, but these tend to be in the mountainous North, which does not have as many hot days, and the tropical South, which has less variability in temperature. We take the average coefficient for $\text{Temperature}^{2}$ across all countries in Fig. 4C. Note that we need to rescale the trip distribution in each country for the effect sizes to be comparable, so the value of each coefficient is difficult to interpret (in Fig. S7, we plot the curves from each ARIMAX formula to show how $\text{Temperature}^{2}$ influences the prediction). We filter out geohashes according to the length of the time series, because very short series could bias the results; generally, as this threshold of minimum observations rises, so too does the effect size.
@@
Finally, we identify the optimal temperature by making predictions across the spectrum for each geohash3 and observing where the predictions begin to fall. According to these predictions, the mean and median optimal temperatures for India are $31.1^{\circ}$ and $31.8^{\circ}$—in line with our estimates from the GAM. (These mirror what we see in S7, which plots the prediction curves for each geohash3 ARIMAX.) Although not a perfect match with the above findings, these results strengthen our assertion that extreme heat can limit activity. We document a noisy relationship between temperature optima across geohashes and the mean temperature of those geohashes ($\rho=0.25$ across all countries, $p<0.05$) in Fig. 4D: as mean temperature increases, the optimal temperature for activity also increases. This provides further evidence that populations adapt to climatic conditions.
@@
We next turn to models of Earth’s climate [51] to make estimates of how activity will be impacted in future. In particular we use data from CMIP6 simulations, extracting the estimates from the Max Planck Institute because these estimates show the best fit to past trends [52]. We substitute each 2019 temperature-day value with its 2050 counterpart from these climate models. We make these projections knowing that our own study suggests that optimal temperatures are often related to the climate of the area, and thus that the population of a given area may be able to adapt to local temperatures. Nevertheless rapid deviation from these temperatures as global warming accelerates [9] might still be a shock to communities who have acclimated to past temperatures. Furthermore, our data suggest that poor areas suffer the worst effects, and these areas might be the least able to cope with changes to the local climate going forward. As a modeling exercise, this provides us with a worst case, but a plausible one at that.
@@
CMIP6 projections (5A) indicate 100–150 annual days above $40^{\circ}$ in Northern states (e.g., Rajasthan), and (5B) suggests 25 days above $45^{\circ}$. We distinguish between large urban areas (top 10 FUAs by population) and small cities (remaining FUAs). On any given day in June when such temperatures are most likely, not a weekend nor a holiday, our model estimates that a single day at $40^{\circ}$ temperatures to reduce activity in large cities areas by 2% and in small ones by 6%, shown in Fig. 5C. Because megacities generate more activity, they experience a small relative change but a large absolute loss. As we see in Fig. 5D, these losses grow to 4% and 8% for large and small urban areas, respectively, during a single day of temperatures at $45^{\circ}$.
@@
We explore the same projections in Mexico, which show the same strong effects from extreme heat and mild effects from temperatures that will occur often. An important difference to note is that while the Southern, tropical areas in Mexico will have the highest average temperature, with Fig.6A showing most days over $35^{\circ}$ by 2050, but the desert near the Northern border—where much of the country’s manufacturing happens—will experience more heat waves, highlighted in Fig. 6B. Because the literature indicates some adaptation to heat, these heat waves in the North could disrupt activity more than stably higher temperatures in the South; although heat waves will be less common in temperate regions, these areas are most at risk of disruption when heat waves do occur.
In Fig. S8, we show similar results for Indonesia, with many days over $35^{\circ}$ in Sumatra and Borneo. While temperatures of $40^{\circ}$ will be rarer, shown in B they are forecast to happen despite never appearing in our 2019 data. Because Indonesia is in the tropics and does not experience much temperature variability, our model projects strong effects for heat that exists beyond the current distribution. Indonesia is planning to move its capital to the island of Borneo, from Java, by 2045. While this megaproject intends to curb the effects of subsidence, temperatures in this new region, according to CMIP6 data, will experience more hot days than in the existing capital region. Shown in Fig. S8C and D, the consequences of these temperatures, according to our modeling, are similar to the effects we see in India: small cities areas with larger relative effects and larger urban areas with larger aggregate disruptions.